OpenStudy (anonymous):

convert each angle in degrees to radians. Express your answer as a multiple of pie. 120 degrees

OpenStudy (anonymous):

360 degrees = 2 pi 120 degrees = ? pi

OpenStudy (lgbasallote):

just multiply the degrees by pi/180 \[\huge 120^o \implies 120 \times \frac{\pi}{180}\]

OpenStudy (anonymous):

@lgbasallote I keep getting the answer wrong.

OpenStudy (lgbasallote):

why so? \[\huge 120 \times \frac{\pi}{180} \implies \frac{120}{180} \times \pi \implies \frac 23 \times \pi \implies \frac{2\pi}{3}\] is that what you're getting?

OpenStudy (anonymous):

@lgbasallote can you explain what you did to get 2/3. I know you multiply but how 120/180xpie = 2/3

OpenStudy (lgbasallote):

\[\huge \frac{120}{180} \implies \frac{60 \times 2}{60 \times 3} \implies \frac{\cancel{60} \times 2}{\cancel{60} \times 3} \implies \frac 23\]